# Extending wavelet regularity beyond Gevrey classes

### v1

@@ 2.1 An example of the extended Gevrey function
First, we discuss the complex extension of the functions defined in (3). For our purposes, it is sufficient to consider unbounded sectors $S_{\gamma}$ given in (13) with $0<\gamma\leq 2$, taking into account only the principal branch of the logarithm.
If $T$ and $S$ are two unbounded sectors, we say that $T$ is a proper subsector of $S$ if $\overline{T}\subseteq S\cup\{0\}$. By taking $\mathcal{R}$ as our initial set, we ensure that the vertex 0 is not included in T{T}. For example, note that $S_{2}=\mathbb{C}\setminus(-\infty,0]$ and $S_{1}=\{z\in\mathbb{C}\mid\mathrm{Re}(z)>0\}$; in this case, $S_{1}$ is a proper subsector of $S_{2}$.
@@
    For every proper subsector $T$ in $S_{2}$ there exist $A_{1},A_{2},\rho^{\prime},\rho^{\prime\prime}>0$ such that
@@
$a)$ Note that $g_{\sigma}=g_{\tau_{\sigma},\sigma}$, where $g_{\tau,\sigma}$ is the function introduced in Lemma 2.15. Therefore, $g_{\sigma}$ is positive and strictly increasing. Moreover, [23, Theorem 3.1] implies that $g_{\sigma}(k)=\omega_{\sigma}(\ln(1+|k|))$, $k>0$, is equivalent to a BMT function (see also Example 1.3). Hence it satisfies the BMT condition $(\alpha)$, and then (15) follows by successive application of this condition.
@@
The fact that $f_{\rho,\sigma}$ is a decreasing function in terms of $\rho$, it follows directly from the definition of $f_{\rho,\sigma}$ in (3). For a fixed $x\in\mathbb{R}$, we note
@@
$C_{2}f_{{\rho_{\sigma}},\sigma}\left(K_{2}/|z|\right)\leq|f_{{\rho_{\sigma}},\sigma}\left(1/z\right)|\leq C_{2}f_{{\rho_{\sigma}},\sigma}\left(K_{1}/|z|\right),\quad z\in T.$ (17)
@@
Proof of the part Lemma 2.17 is technical and therefore given in the Appendix. In the following proposition we capture the regularity of $f_{\rho_{\sigma},\sigma}$ by means of the classes $\mathcal{E}_{\sigma}(\mathbb{R})$.
Proposition 2.18.
For $\sigma>1$ and $\rho_{\sigma}=e^{-1/\sigma}$, let $f_{\rho_{\sigma},\sigma}$ be as in (18). Then it holds $f_{\rho_{\sigma},\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$.
@@
Note that $|z-x|=\frac{x}{2}$ implies $\frac{x}{2}<|z|\leq\frac{3x}{2}$. Using Cauchy’s integral formula and (16) we obtain
@@
$\leq\frac{j!2^{j}}{\pi x^{j+1}}\int_{|z-x|=\frac{x}{2}}c_{1}\,f_{\rho^{\prime},\sigma}\left(K|z|\right)\,|dz|$
$\leq\frac{C_{1}j!2^{j}}{\pi x^{j+1}}\int_{|z-x|=\frac{x}{2}}f_{\rho^{\prime},\sigma}\left(\frac{3Kx}{2}\right)\,|dz|$
$=\frac{C_{1}j!2^{j}}{x^{j}}\,f_{\rho^{\prime},\sigma}\left(\frac{3Kx}{2}\right),\quad x>0,\,\,j\in\mathbb{N}_{0},$ (19)
for suitable constants $C_{1},\rho^{\prime},K>0$, where we also used that $f_{\rho,\sigma}$ are increasing.
@@
$f_{\rho^{\prime},\sigma}\left(\frac{3Kx}{2}\right)=\exp\left\{-\rho^{\prime}g_{\sigma}\Big(\frac{2}{3K}\frac{1}{x}\Big)\right\}\leq L\exp\left\{-\rho^{\prime\prime}g_{\sigma}\Big(\frac{1}{x}\Big)\right\}=Lf_{\rho^{\prime\prime},\sigma}\left(x\right),$ (20)
@@
$\left|\frac{d^{j}}{dt^{j}}f_{\rho_{\sigma},\sigma}(x)\right|\leq C_{1}(2C)^{j^{\sigma}}\frac{f_{\rho^{\prime\prime},\sigma}\left(x\right)}{x^{j}}\leq C^{j^{\sigma}+1}_{2}M^{\tau,\sigma}_{j},\quad x\in\mathbb{K},\;j\in\mathbb{N}_{0},$
@@
Moreover, since the function $z^{\alpha}$, $\alpha>0$, is analytic on $S_{2}=\mathbb{C}\setminus(-\infty,0]$, part $v)$ of Proposition 2.5 implies that $f_{\rho,\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$, and therefore Proposition 2.18 holds as well.
@@
$f_{\rho,\sigma}(t)=\int_{0}^{t}\frac{(t-x)^{p-1}}{(p-1)!}f^{(p)}_{\rho,\sigma}(x)\,dx\leq\frac{t^{p}}{p!}\,C^{p^{\sigma^{\prime}}+1}M^{\tau_{0},\sigma^{\prime}}_{p}\leq C^{\prime}\,t^{p}\,M^{\tau,\sigma}_{p},\qquad p\in\mathbb{N},$
for arbitrary $\tau>0$ and some constant $C^{\prime}>0$ (depending on $\tau$). This contradicts the left-hand side of (10).
@@
Note that Propositions 2.5 (part (i)) and 2.19 imply that, for arbitrary $\rho>0$ and $\sigma>1$, holds
$f_{\rho,\sigma}\notin\bigcup_{t>1}\mathcal{G}_{t},$
where $\mathcal{G}_{t}$ denotes the Gevrey classes.
@@ 3 Wavelet construction
Recall that ${\rm coz}\,m_{0}$ denotes the cozero set (complement of a zero set) of $m_{0}$. Let us start with the following results from [7, 8].
@@
Note that $\frac{2\pi}{3}\notin\operatorname{int}(\operatorname{supp}m_{0})$ suggests that $m_{0}$ vanishes near $\frac{2\pi}{3}$; in our construction, however, we control its local decay, and obtain a non–band-limited scaling function $\varphi$.
If $\operatorname{coz}m_{0}\cap[0,\pi]\subset\left[0,\frac{4}{5}\pi\right)$, note that property (iii) in (21) implies
@@
Moreover, $|m_{0}(\xi)|\leq 1$ for all $\xi\in\mathbb{R}$, and since $m_{0}\big(-\frac{2\pi}{3}\big)=0$, we also have $|m_{0}\left(\frac{\pi}{3}\right)|=1$.
In the following theorem, we control the local behavior of $m_{0}(\xi)$ using the functions $f_{\rho,\sigma}$ defined in (3), to obtain a global decay estimates for $\hat{\varphi}$.
@@
Let $\sigma>1$ and let $m_{0}$ be a low-pass filter that satisfies conditions $i)$, ii)ii) and $iii)$ in (21) such that $\operatorname{coz}\,m_{0}\cap[0,\pi]\subset[0,\frac{4\pi}{5})$ and $\varphi$ be an associated scaling function. Moreover, let functions $f_{\rho,\sigma}$ be as in (3) and ${{\mathbf{\Gamma}}}_{\sigma}$ as in Definition 2.10.
If there exist constants $\varepsilon,C_{0},C_{1},\rho_{1},\rho_{0}>0$ such that
@@
where $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}\in(1,\sigma_{0})$ and $\sigma_{0}=\min\{2,\sigma\}$. Consequently, $\varphi\in{{\mathcal{E}}}_{\sigma}\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{{\mathcal{E}}}_{\sigma^{\prime}}$.
@@
where in the second inequality we used the fact that $|m_{0}(\xi)|\leq 1,\xi\in\mathbb{R}$.
@@
$=C_{0}f_{\rho_{0},\sigma}\left(\frac{|\xi-\frac{2^{n+2}}{3}\pi|}{2^{j}}\right)\leq C_{0}f_{\rho_{0},\sigma}\left(\frac{\varepsilon}{2^{j-n_{0}}}\right)\quad,n_{0}\leq j\leq n+1,$ (27)
for $\xi\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi]$, where for the last inequality we used (26) and the fact that $f_{\rho,\sigma}$ is even, and increasing on $\mathbb{R}_{+}$ (see part $a)$ of the Lemma 2.16).
Moreover, since $f_{\rho_{0},\sigma}\left(\frac{\varepsilon}{2^{j-n_{0}}}\right)=\exp\Big\{-\rho_{0}g_{\sigma}\left(\frac{2^{j-n_{0}}}{\varepsilon}\right)\Big\}$, using again part $a)$ of the Lemma 2.16, for $\xi\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi]$ we obtain
@@
It remains to show that for any $\eta>1$, $\widehat{\varphi}\notin\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{\mathbf{\Gamma}}_{\sigma^{\prime}}$, where $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}\in(1,\sigma_{0})$.
@@
$\xi_{n}=\frac{2^{n+2}}{3}\pi-(-1)^{n}\varepsilon\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi],\quad n\in\mathbb{N},\quad\varepsilon<\frac{8\pi}{15}.$ (29)
@@
for $\tilde{C}_{1}=\tilde{C}C_{1}^{-\log_{2}\pi},\nu_{1}=\log_{2}C_{1}>0$.
Take arbitrary $\eta>1$ set $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}\in(1,\sigma_{0})$. Moreover, if we denote $a_{j}=\ln\left(1+\frac{2^{j}}{\varepsilon}\right)$, then $g_{\sigma}\left(\frac{2^{j}}{\varepsilon}\right)=\omega_{\sigma}(a_{j})$, where $\omega_{\sigma}$ is given in (3). Hence using property $(W2)$ of the Lambert $W$ function, part $a)$ of Lemma 2.16, and that $2^{n}\pi\leq\xi_{n}\leq 2^{n+1}\pi$ we obtain
@@
where $C=1/\ln^{\eta}2$. Clearly series $\sum\frac{\ln^{\eta}j}{j^{\eta}}$ is convergent and hence $C_{\eta}<\infty$.
@@
$|\widehat{\varphi}(\xi_{n})|\geq\tilde{C_{1}}|\xi_{n}|^{\nu_{1}}\exp\{-\rho^{\prime}_{1}\,g_{\sigma_{\eta}}(\xi_{n})\}\geq\tilde{C}_{2}\exp\{-\rho^{\prime}_{2}\,g_{\sigma_{\eta}}(\xi_{n})\},\,n\geq n_{1},$ (33)
@@
Take arbitrary $\rho>0$ and $1<\sigma^{\prime}<\sigma_{\eta}$. Then we can write $\frac{1}{1-\sigma^{\prime}}=\frac{1}{1-\sigma_{\eta}}+c_{\sigma^{\prime}}$ for suitable $c_{\sigma^{\prime}}>0$. Moreover, we can choose large $|\xi_{0}|$ such that for $e^{-c_{\sigma^{\prime}}W(\ln(1+|\xi|))}<\rho$ when $|\xi|>|\xi_{0}|$. This is due to the fact that $\lim_{\xi\to\infty}e^{-c_{\sigma^{\prime}}W(\ln(1+|\xi|))}=0$.
Therefore, using the property $(W2)$ we obtain
@@
Choose $n_{3}\in\mathbb{N}$ such that $|\xi_{n}|>|\xi_{0}|$ for $n\geq n_{2}$. Now (30) follows from (33), and (3) for $n^{\prime}_{0}=\max\{n_{1},n_{2}\}$.
The fact that $\varphi\in{{\mathcal{E}}}_{\sigma}\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{{\mathcal{E}}}_{\sigma^{\prime}}$ now follows from Theorem 2.11 (see also Remark 2.12). ∎
@@
Finally we can prove Theorem 1.2 as our main result.
Proof of Theorem 1.2.
It is sufficient to construct a low-pass filter $m_{0}\in\mathcal{E}_{\sigma}\setminus\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}$ that satisfies the assumptions of Theorem 3.3.
@@
$\gamma_{\sigma}(\eta):=\begin{cases}f_{\sigma}(\eta)\quad&,\eta>0\\ 0\quad&,\eta\leq 0,\end{cases}$
@@
$\delta_{\sigma}(\xi):=\begin{cases}\left(\int_{0}^{1}\gamma_{\sigma}(\eta)\gamma_{\sigma}(1-\eta)\,d\eta\right)^{-1}\int_{0}^{\xi}\gamma_{\sigma}(\eta)\gamma_{\sigma}(1-\eta)\,d\eta\quad&,\xi>0\\ 0\quad&,\xi\leq 0.\end{cases}$
It is clear that $\delta_{\sigma}(\xi)=1$ for all $\xi\geq 1$, and note $\delta_{\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$ (this follows from the property $\widetilde{(M.2)^{\prime}}$ of the $M^{\tau,\sigma}_{p}$). By choosing sufficiently small $\varepsilon>0$ and $0<\rho,\rho^{\prime}<1$, we have
$\gamma_{\sigma}(\eta)$ $\leq f_{\rho,\sigma}(\xi)f_{1-\rho,\sigma}(\eta),$
$\gamma_{\sigma}(1-\eta)$ $\geq f_{\rho^{\prime},\sigma}(\xi)f_{1-\rho^{\prime},\sigma}(1-\eta),\quad\eta\leq\xi<\varepsilon.$ (35)
This follows from the fact that $1-\eta\geq 1-\xi\geq\xi$ when $\xi\in[0,\varepsilon)$, and because $f_{\rho,\sigma}$ are increasing (see part $a)$ of the Lemma (2.16)).
Since $\int_{0}^{\varepsilon}f_{\rho,\sigma}(\eta)d\eta<\infty$, for all $\rho>0$ (see Lemma 2.16 c)), note that (3) implies that there exist $\rho_{1},\rho^{\prime}_{1},C,C^{\prime}>0$ such that for all $\xi\in[0,\varepsilon)$ it holds
@@
The desired low-pass filter $m_{0}$ is given by
$m_{0}(\xi)=\sin\left(\frac{\pi}{2}\theta_{\sigma}(\xi)\right),\,\xi\in\mathbb{R}.$ (37)
Indeed, note that Propositions 2.5 (part $iii)$ and part $v)$) and 2.19 imply that $m_{0}\in\mathcal{E}_{\sigma}\setminus\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}$.
Let us prove that $m_{0}$ in (37) satisfies the assumptions of Theorem 3.3.
For all $\xi\in\mathbb{R}$ it holds
@@
Note that the functions $m_{0}(\xi)$ and $\theta_{\sigma}(\xi)$ have the same decay rate near $\xi=\frac{2\pi}{3}$ , because $\lim_{\xi\to\frac{2\pi}{3}}\frac{m_{0}(\xi)}{\theta_{\sigma}(\xi)}=\frac{\pi}{2}$. Moreover, $\theta_{\sigma}(\xi)$ satisfies the inequalities (24). This follows since the term $\left(1-\delta_{\sigma}\left(\frac{5\xi-\pi}{3\pi}\right)\right)$ in (36) is bounded on $\mathbb{R}$, and it is strictly positive in the neighborhood of $\xi_{0}=\frac{2\pi}{3}$. Therefore, $m_{0}(\xi)$ also satisfies (24) and we conclude the proof. ∎

### v2

@@ 2.1 An example of the extended Gevrey function
First, we discuss the complex extension to sectors of the functions defined in (3). For our purposes, it is sufficient to consider unbounded sectors $S_{\gamma}$ given in (13) with $0<\gamma\leq 2$, taking into account only the principal branch of the logarithm.
If $T$ and $S$ are two unbounded sectors, we say that $T$ is a proper subsector of $S$ if $\overline{T}\subseteq S\cup\{0\}$. By taking the Riemann surface of the logarithm $\mathcal{R}$ as our initial set, we ensure that the vertex 0 is not included in T{T}. For example, note that $S_{2}=\mathbb{C}\setminus(-\infty,0]$ and $S_{1}=\{z\in\mathbb{C}\mid\mathrm{Re}(z)>0\}$; in this case, $S_{1}$ is a proper subsector of $S_{2}$.
@@
    For every proper subsector $T$ in $S_{2}$ there exist $C_{1},C_{2},\rho^{\prime},\rho^{\prime\prime}>0$ such that
@@
$a)$ Note that $g_{\sigma}=g_{\tau_{\sigma},\sigma}$, where $g_{\tau,\sigma}$ is the function introduced in Lemma 2.14. Therefore, $g_{\sigma}$ is positive and strictly increasing. Moreover, [25, Theorem 3.1] implies that $g_{\sigma}(k)=\omega_{\sigma}(\ln(1+|k|))$, $k>0$, is equivalent to a BMT function (see also Example 1.2). Hence it satisfies the BMT condition $(\alpha)$, which implies  (15).
@@
The fact that $f_{\rho,\sigma}$ is a decreasing function with respect to $\rho$ follows directly from the definition of $f_{\rho,\sigma}$ in (3). For a fixed $x\in\mathbb{R}$, we note
@@
$C_{2}f_{{\rho_{\sigma}},\sigma}\left(K_{2}/|z|\right)\leq|f_{{\rho_{\sigma}},\sigma}\left(1/z\right)|\leq C_{1}f_{{\rho_{\sigma}},\sigma}\left(K_{1}/|z|\right),\quad z\in T.$ (17)
@@
Proof of the Lemma 2.16 is rather technical and therefore given in the Appendix. In the following proposition we capture the regularity of $f_{\rho_{\sigma},\sigma}$ by means of the classes $\mathcal{E}_{\sigma}(\mathbb{R})$.
Theorem 2.17.
Let $\sigma>1$, $\rho_{\sigma}=e^{-1/\sigma}$, and let $f_{\rho_{\sigma},\sigma}$ be given by (18). Then $f_{\rho_{\sigma},\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$.
@@
Note that $|z-x|=\frac{x}{2}$ implies $\frac{x}{2}\leq|z|\leq\frac{3x}{2}$. Using Cauchy’s integral formula and (16) we obtain
@@
$\leq\frac{j!2^{j}}{\pi x^{j+1}}\int_{|z-x|=\frac{x}{2}}C_{1}\,f_{\rho^{\prime},\sigma}\left(|z|\right)\,|dz|$
$\leq\frac{C_{1}j!2^{j}}{\pi x^{j+1}}\int_{|z-x|=\frac{x}{2}}f_{\rho^{\prime},\sigma}\left(\frac{3x}{2}\right)\,|dz|$
$=\frac{C_{1}j!2^{j}}{x^{j}}\,f_{\rho^{\prime},\sigma}\left(\frac{3x}{2}\right),\quad x>0,\,\,j\in\mathbb{N}_{0},$ (19)
for suitable constants $C_{1},\rho>0$, where we also used that $f_{\rho,\sigma}$ are increasing.
@@
$f_{\rho^{\prime},\sigma}\left(\frac{3x}{2}\right)=\exp\left\{-\rho^{\prime}g_{\sigma}\Big(\frac{2}{3x}\Big)\right\}\leq L\exp\left\{-\rho^{\prime\prime}g_{\sigma}\Big(\frac{1}{x}\Big)\right\}=Lf_{\rho^{\prime\prime},\sigma}\left(x\right),$ (20)
@@
$\left|\frac{d^{j}}{dx^{j}}f_{\rho_{\sigma},\sigma}(x)\right|\leq C_{1}(2C)^{j^{\sigma}}\frac{f_{\rho^{\prime\prime},\sigma}\left(x\right)}{x^{j}}\leq C^{j^{\sigma}+1}_{2}M^{\tau,\sigma}_{j},\quad x\in\mathbb{K},\;j\in\mathbb{N}_{0},$
@@
Moreover, since the function $z^{\alpha}$, $\alpha>0$, is analytic on $S_{2}=\mathbb{C}\setminus(-\infty,0]$, Proposition 2.5 $e)$ implies that $f_{\rho,\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$, and therefore Proposition 2.17 holds as well.
@@
$f_{\rho,\sigma}(x)=\int_{0}^{x}\frac{(x-s)^{p-1}}{(p-1)!}f^{(p)}_{\rho,\sigma}(s)\,ds\leq\frac{x^{p}}{p!}\,C^{p^{\sigma^{\prime}}+1}M^{\tau_{0},\sigma^{\prime}}_{p}\leq C^{\prime}\,x^{p}\,M^{\tau,\sigma}_{p},\qquad p\in\mathbb{N},$
for arbitrary $\tau>0$ and some constant $C^{\prime}>0$ (depending on $\tau$). This contradicts the left-hand side of (10). ∎
@@
Note that Proposition 2.5 $a)$ and Proposition 2.18 imply that, for arbitrary $\rho>0$ and $\sigma>1$, holds $f_{\rho,\sigma}\notin\bigcup_{t>1}\mathcal{G}_{t}$, where $\mathcal{G}_{t}$ denotes the Gevrey classes.
@@ 3 Wavelet construction
Recall that ${\rm coz}\,m_{0}$ denotes the cozero set (complement of a zero set) of $m_{0}$. Let us start with the following results, see [8, Lemmas 2.3 and 3.1] and [9, Theorem 1].
@@
Note that the condition
$\frac{2\pi}{3}\notin\operatorname{int}(\operatorname{supp}m_{0})=\operatorname{int}\big(\overline{\operatorname{coz}m_{0}}\big),$
implies that $m_{0}$ vanishes in a neighborhood of $\frac{2\pi}{3}$. In our construction, we prescribe the local decay of $m_{0}$ near $\frac{2\pi}{3}$, so that
$\frac{2\pi}{3}\notin\operatorname{coz}m_{0},\qquad\text{but}\qquad\frac{2\pi}{3}\in\operatorname{int}(\operatorname{supp}m_{0}).$
If $\operatorname{coz}m_{0}\cap[0,\pi]\subset\left[0,\frac{4}{5}\pi\right)$, note that property $iii)$ in (21) implies
@@
Moreover, $|m_{0}(\xi)|\leq 1$ for all $\xi\in\mathbb{R}$, and since $m_{0}\big(-\frac{2\pi}{3}\big)=0$, we also have $\Big|m_{0}\left(\frac{\pi}{3}\right)\Big|=1$. This yields to non-band-limited scaling function $\varphi$ given in (1).
In the following theorem, we control the local behavior of $m_{0}(\xi)$ using the functions $f_{\rho,\sigma}$ defined in (3), to obtain a global decay estimates for $\hat{\varphi}$. Moreover, we consider $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}$, $\sigma>1$, $\eta>1$. Note that $\sigma_{\eta}\in(1,\sigma_{0})$ where $\sigma_{0}=\min\{2,\sigma\}$.
@@
Let $m_{0}$ be a low-pass filter that satisfies conditions $i)$, ii)ii) and $iii)$ in (21) such that $\operatorname{coz}\,m_{0}\cap[0,\pi]\subset[0,\frac{4\pi}{5})$ and $\varphi$ be the associated scaling function. Moreover, let $\rho>0$, $\sigma>1$, $f_{\rho,\sigma}$ be as in (3) and ${{\mathbf{\Gamma}}}_{\sigma}$ as in Definition 2.10.
Then the following is true: if $m_{0}$ satisfies
@@
for each $\eta>1$. Consequently, $\varphi\in{{\mathcal{E}}}_{\sigma}\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{{\mathcal{E}}}_{\sigma^{\prime}}$.
@@
where in the second inequality we used the fact that $|m_{0}(\xi)|\leq 1,$ $\xi\in\mathbb{R}$, see Remark 3.2.
@@
$=C_{0}f_{\rho_{0},\sigma}\left(\frac{|\xi-\frac{2^{n+2}}{3}\pi|}{2^{j}}\right)\leq C_{0}f_{\rho_{0},\sigma}\left(\frac{\varepsilon}{2^{j-n_{0}}}\right),\quad n_{0}\leq j\leq n+1,$ (27)
for $\xi\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi]$, where for the last inequality we used (26) and the fact that $f_{\rho,\sigma}$ is even, and increasing on $\mathbb{R}_{+}$ (see Lemma 2.15 $a)$).
Moreover, since $f_{\rho_{0},\sigma}\left(\frac{\varepsilon}{2^{j-n_{0}}}\right)=\exp\Big\{-\rho_{0}g_{\sigma}\left(\frac{2^{j-n_{0}}}{\varepsilon}\right)\Big\}$, using again Lemma 2.15 $a)$, for $\xi\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi]$ we obtain
@@
It remains to show that for any $\eta>1$, $\widehat{\varphi}\notin\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{\mathbf{\Gamma}}_{\sigma^{\prime}}$, where $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}$.
@@
$\xi_{n}:=\frac{2^{n+2}}{3}\pi-(-1)^{n}\varepsilon\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi],\quad n\in\mathbb{N},\quad\varepsilon<\frac{8\pi}{15}.$ (29)
@@
for $\tilde{C}_{1}=\tilde{C}C_{1}^{-\log_{2}\pi}$ and $\nu_{1}=\log_{2}C_{1}$.
Take arbitrary $\eta>1$ set $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}\in(1,\sigma_{0})$. Moreover, if we denote $a_{j}=\ln\left(1+\frac{2^{j}}{\varepsilon}\right)$, then $g_{\sigma}\left(\frac{2^{j}}{\varepsilon}\right)=\omega_{\sigma}(a_{j})$, where $\omega_{\sigma}$ is given in (3). Hence using property $(W2)$ of the Lambert $W$ function, Lemma 2.15 $a)$, and that $2^{n}\pi\leq\xi_{n}\leq 2^{n+1}\pi$ we obtain
@@
where $C=1/\ln^{\eta}2$. Clearly series $\sum\frac{\ln^{\eta}j}{j^{\eta}}$ is convergent for $\eta>1$ and hence $C_{\eta}<\infty$.
@@
$|\widehat{\varphi}(\xi_{n})|\geq\tilde{C_{1}}|\xi_{n}|^{\nu_{1}}\exp\{-\rho^{\prime}_{1}\,g_{\sigma_{\eta}}(\xi_{n})\}\geq\tilde{C}_{2}\exp\{-\rho^{\prime}_{2}\,g_{\sigma_{\eta}}(\xi_{n})\},\quad n\geq n_{1},$ (33)
@@
Take arbitrary $\rho>0$ and $1<\sigma^{\prime}<\sigma_{\eta}$. Then we can write
$\frac{1}{1-\sigma^{\prime}}=\frac{1}{1-\sigma_{\eta}}+c_{\sigma^{\prime}},$
for suitable $c_{\sigma^{\prime}}>0$. Moreover, since
$\lim_{\xi\to\infty}e^{-c_{\sigma^{\prime}}W(\ln(1+|\xi|))}=0,$
we can choose $|\xi_{0}|$ large enough to obtain $e^{-c_{\sigma^{\prime}}W(\ln(1+|\xi|))}<\rho$ when $|\xi|>|\xi_{0}|$. Therefore, using the property $(W2)$ we have
@@
Choose $n_{2}\in\mathbb{N}$ such that $|\xi_{n}|>|\xi_{0}|$ for $n\geq n_{2}$, where $\xi_{n}$ are given in (29). Now (30) follows from (33), and (3) for $n^{\prime}_{0}=\max\{n_{1},n_{2}\}$.
The fact that $\varphi\in{{\mathcal{E}}}_{\sigma}\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{{\mathcal{E}}}_{\sigma^{\prime}}$ now follows from Theorem 2.11. ∎
@@
Now we are ready to prove our main result.
Theorem 3.5.
For given $\sigma>1$ there exists an orthonormal wavelet $\psi$ such that
$\psi\in\mathcal{E}_{\sigma}(\mathbb{R})\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}\mathcal{E}_{\sigma^{\prime}}(\mathbb{R})\quad{\rm and}\quad\displaystyle\hat{\psi}\in\mathcal{E}_{\sigma}(\mathbb{R})\backslash\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}(\mathbb{R}),$
for each $\eta>1$ where $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}$.
Proof.
It is sufficient to construct a low-pass filter $m_{0}\in\mathcal{E}_{\sigma}\setminus\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}$ that satisfies the assumptions of Theorem 3.3. The decay and regularity of $\hat{\psi}$ then follow from (2).
@@
$\gamma_{\sigma}(\mu):=\begin{cases}f_{\sigma}(\mu),\quad&\mu>0\\ 0,\quad&\mu\leq 0,\end{cases}$
@@
$\delta_{\sigma}(\xi):=\begin{cases}\left(\int_{0}^{1}\gamma_{\sigma}(\mu)\gamma_{\sigma}(1-\mu)\,d\mu\right)^{-1}\int_{0}^{\xi}\gamma_{\sigma}(\mu)\gamma_{\sigma}(1-\mu)\,d\mu,\quad&\xi>0\\ 0,\quad&\xi\leq 0.\end{cases}$
It is clear that $\delta_{\sigma}(\xi)=1$ for all $\xi\geq 1$, and note $\delta_{\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$. This follows from the property $\widetilde{(M.2)^{\prime}}$ of the $M^{\tau,\sigma}_{p}$, and Proposition 2.5 $c)$. By choosing sufficiently small $\varepsilon>0$ and $0<\rho,\rho^{\prime}<1$, we have
$\gamma_{\sigma}(\mu)$ $\leq f_{\rho,\sigma}(\xi)f_{1-\rho,\sigma}(\mu),$
$\gamma_{\sigma}(1-\mu)$ $\geq f_{\rho^{\prime},\sigma}(\xi)f_{1-\rho^{\prime},\sigma}(1-\mu),\quad\mu\leq\xi<\varepsilon.$ (35)
This follows from the fact that $1-\mu\geq 1-\xi\geq\xi$ when $\xi\in[0,\varepsilon)$, and because $f_{\rho,\sigma}$ are increasing (see Lemma (2.15) $a)$).
Since $\int_{0}^{\varepsilon}f_{\rho,\sigma}(\mu)d\mu<\infty$, for all $\rho>0$ (see Lemma 2.15 $c)$), the estimates in (3) imply that there exist $\rho_{1},\rho^{\prime}_{1},C,C^{\prime}>0$ such that
@@
The desired low-pass filter $m_{0}$ is then given by
$m_{0}(\xi)=\sin\left(\frac{\pi}{2}\theta_{\sigma}(\xi)\right),\quad\xi\in\mathbb{R}.$ (37)
Indeed, note that Proposition 2.5 $c)$ and $e)$ applied to $\mathcal{E}_{\sigma}$, and Proposition 2.18 imply that $m_{0}\in\mathcal{E}_{\sigma}\setminus\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}$.
Let us prove that $m_{0}$ in (37) satisfies the assumptions of Theorem 3.3. For all $\xi\in\mathbb{R}$ it holds
@@
Note that the functions $m_{0}(\xi)$ and $\theta_{\sigma}(\xi)$ have the same decay rate near $\xi=\frac{2\pi}{3}$ , because $\lim_{\xi\to\frac{2\pi}{3}}\frac{m_{0}(\xi)}{\theta_{\sigma}(\xi)}=\frac{\pi}{2}$. Moreover, $\theta_{\sigma}(\xi)$ satisfies the inequalities (24). This follows since the term $\left(1-\delta_{\sigma}\left(\frac{5\xi-\pi}{3\pi}\right)\right)$ in (36) is bounded on $\mathbb{R}$, and it is strictly positive in the neighborhood of $\xi_{0}=\frac{2\pi}{3}$. Therefore, $m_{0}(\xi)$ also satisfies (24). ∎